# xApp Conflict Mitigation with Scheduler

### v2

@@ 5.2 Training of Scheduler
[…]
5.3 Simulation Results
@@
It is assumed that an intent is received as: ”maximize the total transmission rate across all O-RUs” and the Non-RT RIC selects xApps, $\{\mathcal{X}_{1},\mathcal{X}_{2},\mathcal{X}_{3},\mathcal{X}_{4}\}$ based on target, $f$, of the total transmission rate defined in (25). This set of trained xApps is deployed to xApp IH to be managed by the scheduler.
Fig. 6 compares the performance of multiple active trained xApp variations, including RBG-only$(\mathcal{X}_{4})$, Power-only $(\mathcal{X}_{2})$, and Both ($(\mathcal{X}_{2})$, $(\mathcal{X}_{4})$) deployed independently; Both ($(\mathcal{X}_{2})$, $(\mathcal{X}_{4})$) with scheduler that considers only two A2C-trained xApps while retaining the last action from a deactivated xApp; and a 4 xApps ($(\mathcal{X}_{1})$, $(\mathcal{X}_{2})$, $(\mathcal{X}_{3})$, $(\mathcal{X}_{4})$) with scheduler scenario where the scheduler selects one xApp from each group of power and RBG allocations. These comparisons are conducted with context variables categorized into low, mid, and high values. It can be observed that the scheduler utilizing both trained A2C xApps demonstrates improved performance compared to the conflicting case; however, its performance remains inferior to that achieved through individual deployment of the xApps. Extending the set of available xApps to include baseline xApps and permitting the scheduler to select from this expanded group yields the best results, surpassing even the performance of individual xApp deployments.
6 CONCLUSION
@@
In the O-RAN network considered here, multiple distributed O-RUs rely on xApps for essential tasks, such as power and resource block allocations. Despite its capacity for data-driven control, the O-RAN paradigm necessitates robust coordination among pre-trained xApps, as their actions can occasionally conflict. This work underscores the inherent challenges in estimating KPMs for diverse action combinations of active xApps in the absence of an accurate digital twin, and highlights the context-dependent nature of conflicts in automated network management. Recognizing that pre-trained and validated xApps cannot be modified post-deployment, this study introduces an intent-driven, scheduler-based conflict mitigation framework that selects active xApps based on contextual variables and target KPMs, without necessitating further xApp re-training. Through simulation of an indirect conflict scenario involving Power and RGB allocation xApps, and by leveraging the A2C method for training both the xApps and the scheduler, it is demonstrated that while a scheduler utilizing only A2C-trained xApps improves performance relative to independently deployed conflicting xApps, its performance is suboptimal compared to individual deployments. Notably, extending the available set of xApps to include baseline configurations and enabling the scheduler to choose from this expanded pool yielded the most favorable results. These findings advocate for further exploration into adaptive scheduling mechanisms and a broader xApp selection framework to enhance conflict resolution and ensure optimal network performance in dynamic, real-world deployments. The ideal scheduler should dynamically manipulate pre-trained xApps based on context-specific network state variables and be easily updated to accommodate various intents and combinations of xApps.

### v3

@@ V-B Training of Scheduler
Safety Layer: Confidence‑Gated Fallback Mechanism
Despite extensive training, the A2C scheduler can encounter context vectors that fall outside its training distribution. To bound performance in such regimes, the scheduler can be encapsulated within a lightweight confidence gate: whenever the critic’s value estimate falls below a threshold anomalously, the gate suppresses the RL action and temporarily invokes a pre-determined deterministic fallback policy with a known performance floor.
Let $V_{\phi}(\mathbf{s}_{\dagger})$ denote the critic’s value estimate at scheduling instant $\dagger$ and maintain exponentially weighted estimates of its mean and dispersion with forgetting factor $\beta\in(0,1]$ [121]:
$m_{\dagger}$ $=(1-\beta)\,m_{\dagger-1}+\beta\,V_{\phi}(\mathbf{s}_{\dagger}),$ (25)
$\sigma_{\dagger}$ $=(1-\beta)\,\sigma_{\dagger-1}+\beta\bigl{|}V_{\phi}(\mathbf{s}_{\dagger})-m_{\dagger-1}\bigr{|}.$ (26)
A state is classified as out‑of‑distribution if the zz‑score
$z_{\dagger}=\frac{V_{\phi}(\mathbf{s}_{\dagger})-m_{\dagger}}{\sigma_{\dagger}}$
falls below a threshold defined by MNO. When this condition is met, the learned action $\mathbf{a}_{\dagger}$ is overridden for the next $T_{\text{back}}$ decisions by a deterministic fallback policy $\pi^{\text{safe}}$: (i) equal resource allocation or (ii) the single xApp that achieved the highest offline average reward. Note that the statistics $(m_{\dagger},\sigma_{\dagger})$ are frozen during the back‑off window to avoid bias, and the gate re‑arms automatically when the timer elapses.
V-C Simulation Results of Testing
@@
It is assumed that an intent is received as: ”maximize the total transmission rate across all O-RUs” and the Non-RT RIC selects xApps, $\{\mathcal{X}_{1},\mathcal{X}_{2},\mathcal{X}_{3},\mathcal{X}_{4}\}$ based on target, $f$, of the total transmission rate defined in (24). This set of trained xApps is deployed to xApp IH to be managed by the scheduler. To evaluate the performance of schedulers, tests are explicitly conducted against the conflict scenario, involving the simultaneous deployment of the power and RBG xApps. The evaluations are performed across nine distinct combinations of context variables, specifically $c_{\dagger}^{1}\in\{2,5,8\}\text{ Mbps}$ and $c_{\dagger}^{2}\in\{5,25,45\}\text{ m/s}$, thereby comprehensively capturing diverse system behaviors.
Figure 7: Mean leftover bits at the end of each episode
Fig. 6 compares the performance of multiple active trained xApp variations, including Both ($(\mathcal{X}_{1})$, $(\mathcal{X}_{2})$) deployed independently (i.e. conflicting case); Method 1: Both ($(\mathcal{X}_{1})$, $(\mathcal{X}_{2})$) with scheduler and a Method 2: 4 xApps ($(\mathcal{X}_{1})$, $(\mathcal{X}_{2})$, $(\mathcal{X}_{3})$, $(\mathcal{X}_{4})$) with scheduler. It is apparent that variations in the mean user speed have a lesser impact on performance compared to increases in the mean data arrival rate. Furthermore, it is evident that employing the scheduler that retains the previous action (i.e., Method 1) results in improved performance relative to the conflicting scenario. However, its effectiveness remains limited due to the constrained action set provided by the independently operating A2C-based xApps. Each xApp determines its actions autonomously, without direct influence from the scheduler, as they are individually optimized to maximize their own objectives by tuning different NCPs. Extending the available action space by incorporating baseline xApps (i.e., Method 2) and enabling the scheduler to dynamically choose from this expanded set results in the best overall performance. This framework enhances the scheduler’s adaptability by broadening its action space through the integration of baseline actions, enabling flexible selection between pre-trained xApp policies and baseline options based on the prevailing network context.
Furthermore, Fig. 7 illustrates the mean leftover (i.e., discarded) bits at the end of each episode, arising from insufficient capacity, as defined in (7). It is also evident that the degree of the conflict depends on the context variables, and the implementation of scheduler designs contributes to reducing these conflicts and enhancing overall performance. Consistent with Fig. 6, the scheduler employing baseline policies in addition to A2C xApps (i.e., Method 2) exhibits superior performance compared to the scheduler utilizing exclusively A2C xApps (i.e., Method 1).
VI CONCLUSION
@@
In the O-RAN network considered here, multiple distributed O-RUs rely on xApps for essential tasks, such as power and resource block allocations. Despite its capacity for data-driven control, the O-RAN paradigm necessitates robust coordination among pre-trained xApps, as their actions can occasionally conflict. This work underscores the inherent challenges in estimating KPMs for diverse action combinations of active xApps in the absence of an accurate digital twin, and highlights the context-dependent nature of conflicts in automated network management. Recognizing that pre-trained and validated xApps cannot be modified post-deployment, this study introduces an intent-driven, scheduler-based conflict mitigation framework that selects active xApps based on contextual variables and target KPMs, without necessitating further xApp re-training. Through simulation of an indirect conflict scenario involving Power and RBG allocation xApps, and by leveraging the A2C method for training both the xApps and the scheduler, it is demonstrated that a scheduler utilizing only A2C-trained xApps improves performance relative to independently deployed conflicting xApps. Notably, extending the available set of xApps to include baseline configurations and enabling the scheduler to choose from this expanded pool yielded the most favorable results. These findings advocate for further exploration into adaptive scheduling mechanisms and a broader xApp selection framework to enhance conflict resolution and ensure optimal network performance in practical deployment scenarios through advanced testbed platforms. The ideal scheduler should dynamically manipulate pre-trained xApps based on context-specific network state variables and be easily updated to accommodate various intents and combinations of xApps.